OpenStudy (anonymous):

I'm studying a back test for my Mathematical Analysis course, but I'm not sure how to answer one of the questions. Give an example of a sequence {yn} such that lim (n->Infinity) |yn-y(n+1)|=0 but is not a Cauchy Sequence.

OpenStudy (anonymous):

i don't have a ready answer, but i have an idea

OpenStudy (turingtest):

oh good Q

OpenStudy (anonymous):

the difference between this and a cauchy sequence is that this says adjacent elements of the sequence converge to zero, not arbitrary ones

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

try \(a_n\) as the nth partial sum of the harmonic series

OpenStudy (anonymous):

i think that is right, and it generalizes to \[\lim|a_{n+r}-a_n|=0\]

OpenStudy (anonymous):

for fixed \(r\) of course.

OpenStudy (anonymous):

@Aylin is the answer clear?

OpenStudy (anonymous):

I was just about to come back and ask if y{1}=1 and y{n+1}=y{n}+1/n would work.

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

point is the harmonic series diverges, so you can never find an \(N\) such that if \(m,n>N\) you would have \[\sum_1^n\frac{1}{k}-\sum_1^m\frac{1}{k}<\epsilon\] in fact this difference can be made larger than and larger. but \[\sum_1^n\frac{1}{k}-\sum_1^{n+1}\frac{1}{k}=\frac{1}{n+1}\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

yw